# Improving Online-to-Nonconvex Conversion for Smooth Optimization via Double Optimism

### v2

@@ C.1 Helpful inequalities
This effectively decouples the noise in $\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|$ into separate, more manageable components. We now focus on bounding the expectation of the first term in the preceding inequality. By applying Jensen’s inequality and invoking Assumption 2.3, we obtain: $\operatorname*{\mathbb{E}}\Bigl[2\sqrt{\sum_{n=1}^{T}(\|{\mathbf{g}}_{n}-\bar{{\mathbf{g}}}_{n}\|^{2}+\|{\mathbf{h}}_{n}-\bar{{\mathbf{h}}}_{n}\|^{2})}\Bigr]\allowbreak\leq 2\sqrt{\sum_{n=1}^{T}\operatorname*{\mathbb{E}}[\|{\mathbf{g}}_{n}-\bar{{\mathbf{g}}}_{n}\|^{2}+\|{\mathbf{h}}_{n}-\bar{{\mathbf{h}}}_{n}\|^{2}]}\leq 2\sqrt{2}\sigma\sqrt{T}$. Therefore, it only remains to upper bound
@@
  • $16C_{1}^{\frac{3}{2}}\hat{L}_{1}^{*}\gamma^{\frac{1}{2}}\frac{D}{T}\leq\max\left(16^{\frac{2}{3}}L_{2}^{\frac{1}{3}}(\hat{L}_{1}^{*})^{\frac{2}{3}}\gamma^{\frac{1}{3}}C_{1}D^{\frac{4}{3}},32C_{1}^{\frac{3}{2}}\hat{L}_{1}^{*}\gamma^{\frac{1}{2}}\frac{D}{M}\right)\leq\newline \max\left(7L_{2}^{\frac{1}{3}}(\hat{L}_{1}^{*})^{\frac{2}{3}}\gamma^{\frac{1}{3}}C_{1}D^{\frac{4}{3}},32C_{1}^{\frac{3}{2}}\hat{L}_{1}^{*}\gamma^{\frac{1}{2}}\frac{D}{M}\right)$;
@@
  • $2\frac{F({\mathbf{x}}_{0})-F^{*}}{DM}\leq 2\cdot 10^{\frac{3}{7}}\gamma^{\frac{1}{7}}C_{1}^{\frac{3}{7}}(\hat{L}_{1}^{*})^{\frac{2}{7}}L_{2}^{\frac{1}{7}}\frac{(F({\mathbf{x}}_{0})-F^{*})^{\frac{4}{7}}}{M^{\frac{4}{7}}}+2\cdot 1.5^{\frac{5}{7}}\frac{(F({\mathbf{x}}_{0})-F^{*})^{\frac{2}{7}}}{M^{\frac{2}{7}}}C_{2}^{\frac{4}{7}}L_{2}^{\frac{1}{7}}\sigma^{\frac{4}{7}}\leq\newline 6\gamma^{\frac{1}{7}}C_{1}^{\frac{3}{7}}(\hat{L}_{1}^{*})^{\frac{2}{7}}L_{2}^{\frac{1}{7}}\frac{(F({\mathbf{x}}_{0})-F^{*})^{\frac{4}{7}}}{M^{\frac{4}{7}}}+3\frac{(F({\mathbf{x}}_{0})-F^{*})^{\frac{2}{7}}}{M^{\frac{2}{7}}}C_{2}^{\frac{4}{7}}L_{2}^{\frac{1}{7}}\sigma^{\frac{4}{7}}$;

### v3

@@ C.1 Helpful inequalities
Lemma C.1 appears in [KLBC19, AKC22, LYC18]; for the sake of completeness, we additionally provide its proof here.
@@
This effectively decouples the noise in $\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|$ into separate, more manageable components. We now focus on bounding the expectation of the first term in the preceding inequality. By applying Jensen’s inequality and invoking Assumption 2.3 $\operatorname*{\mathbb{E}}\Bigl[2\sqrt{\sum_{n=1}^{T}(\|{\mathbf{g}}_{n}-\bar{{\mathbf{g}}}_{n}\|^{2}+\|{\mathbf{h}}_{n}-\bar{{\mathbf{h}}}_{n}\|^{2})}\Bigr]\leq 2\sqrt{\sum_{n=1}^{T}\operatorname*{\mathbb{E}}[\|{\mathbf{g}}_{n}-\bar{{\mathbf{g}}}_{n}\|^{2}+\|{\mathbf{h}}_{n}-\bar{{\mathbf{h}}}_{n}\|^{2}]}\leq 2\sqrt{2}\sigma\sqrt{T}$. Therefore, it only remains to upper bound
@@
  • $16C_{1}^{\frac{3}{2}}\hat{L}_{1}^{*}\gamma^{\frac{1}{2}}\frac{D}{T}\leq\max\left(16^{\frac{2}{3}}L_{2}^{\frac{1}{3}}(\hat{L}_{1}^{*})^{\frac{2}{3}}\gamma^{\frac{1}{3}}C_{1}D^{\frac{4}{3}},32C_{1}^{\frac{3}{2}}\hat{L}_{1}^{*}\gamma^{\frac{1}{2}}\frac{D}{M}\right)$
    $\leq\max\left(7L_{2}^{\frac{1}{3}}(\hat{L}_{1}^{*})^{\frac{2}{3}}\gamma^{\frac{1}{3}}C_{1}D^{\frac{4}{3}},32C_{1}^{\frac{3}{2}}\hat{L}_{1}^{*}\gamma^{\frac{1}{2}}\frac{D}{M}\right)$;
@@
  • $2\frac{F({\mathbf{x}}_{0})-F^{*}}{DM}\leq 2\cdot 10^{\frac{3}{7}}\gamma^{\frac{1}{7}}C_{1}^{\frac{3}{7}}(\hat{L}_{1}^{*})^{\frac{2}{7}}L_{2}^{\frac{1}{7}}\frac{(F({\mathbf{x}}_{0})-F^{*})^{\frac{4}{7}}}{M^{\frac{4}{7}}}+2\cdot 1.5^{\frac{5}{7}}\frac{(F({\mathbf{x}}_{0})-F^{*})^{\frac{2}{7}}}{M^{\frac{2}{7}}}C_{2}^{\frac{4}{7}}L_{2}^{\frac{1}{7}}\sigma^{\frac{4}{7}}$
    $\leq 6\gamma^{\frac{1}{7}}C_{1}^{\frac{3}{7}}(\hat{L}_{1}^{*})^{\frac{2}{7}}L_{2}^{\frac{1}{7}}\frac{(F({\mathbf{x}}_{0})-F^{*})^{\frac{4}{7}}}{M^{\frac{4}{7}}}+3\frac{(F({\mathbf{x}}_{0})-F^{*})^{\frac{2}{7}}}{M^{\frac{2}{7}}}C_{2}^{\frac{4}{7}}L_{2}^{\frac{1}{7}}\sigma^{\frac{4}{7}}$;